OpenStudy (anonymous):

a space vehicle accelerates uniformly from 80m/s at t=0 to 164m/s at t=10.0s how far did it move between t=2.0s and t=6.0s?

OpenStudy (badhi):

find the acceleration from a = (v-u)/t then find the velocity at t=2 from v = u + at find the distance from t=2 to t = 6 from s = ut + 1/2 at^2

OpenStudy (anonymous):

may I ask what the "u" stands for

OpenStudy (badhi):

u = velocity at the begining of the considered time period v = velocity at the end of the considered time period

OpenStudy (anonymous):

oh thank you !